# Traversable Wormholes with Spontaneous Symmetry Breaking

### v1

@@ Abstract
We present an exact, spherically symmetric, static solution of the Einstein field equations minimally coupled to a self-interacting scalar field. The solution does not exhibit any zero proper volume singularity at the center, and therefore, curvature scalars are always regular. We study the property of radial null geodesics and derive that the metric can describe either a two-way or a one-way traversable wormhole depending on certain parameter ranges. Notably, it lacks a Schwarzschild limit. The scalar field exhibits spontaneous symmetry breaking within the coordinate range where a wormhole throat forms. It can be seen as a suggestion that spontaneous symmetry breaking may act as a threshold for wormhole throat formation. Additionally, we compute the radius of the photon sphere, the Lyapunov exponent, the shadow radius, and the innermost stable circular orbits for this spacetime metric.
@@ 1 Introduction
Although the General Theory of Relativity (GR) is widely regarded as the most accurate description of gravity to date, the number of exact solutions to its field equations remains relatively small. This limitation is primarily due to the fundamental challenges involved in solving the coupled partial differential equations that form the basis of the theory. In any case, the spherically symmetric field equations of GR are inherently quite restrictive. A clear example of this rigidity is found in the fact that, for a vacuum (where $T_{\mu\nu}=0$), the Schwarzschild solution stands as the unique static and isotropic solution that satisfies the condition of asymptotic flatness [1, 2, 3]. This result is central to the formulation of Jebsen-Birkhoff’s theorem [4, 5]. The Schwarzschild solution represents a singularity with zero proper volume, surrounded by a horizon, and depends on two parameters: $r_{s}$, which indicates the location of the horizon, and $r$, which defines the position of the singularity. When matter is incorporated into the field equations, they become even more restrictive and less predictable. Some partial solutions have been obtained in specific cases, such as GR coupled to self-interacting scalar fields of dilaton or axion types or when coupled to Maxwell electrodynamics or a Yang-Mills field. However, static field configurations in GR are heavily constrained outside the event horizon, as demonstrated by the no-hair theorems.
For a simple minimally coupled self-interacting scalar field $\phi$, the contribution to the energy-momentum tensor is written as
@@
$V(\phi)$ is the self-interaction of the scalar field. The scalar field profile is imposed upon by a requirement that an $r\rightarrow\infty$ limit should generate a constant value of the scalar field $\phi(r)$ (one can choose to normalize this constant value to zero). Alongwith this, a requirement of asymptotic flatness dictates that the self-interaction potential should have a local extremum at the origin of the scalar field. In fact, we can use these conditions to demonstrate the no-hair theorem starting from the scalar field evolution equation
@@
For a static case, only the $g^{rr}$ term contributes to the above equation. We multiply the equation with $\phi(r)$ and do integration by parts from the horizon to infinity, yielding:
@@
Usually, one realizes a horizon $h$ at the largest zero of $g^{rr}$. Thus in principle it is possible to integrate the above equation for all $g^{rr}\geq 0$. If we assume regularity of $\phi$ and $\sqrt{-g}$ at the horizon, then the boundary terms can be dropped (also keeping in mind an asymptotic fall-off of $\phi(r)$ as $r\rightarrow\infty$). In this context, the condition $\phi\frac{dV}{d\phi}\geq 0$ for all $r\geq h$ essentially implies that the self-interaction potential has a local minimum at $\phi=0$ and no local maximum in the $\phi$-range considered outside the horizon. Under this condition, the integrands on both sides of the above equation are clearly non-negative. However, the integral on LHS brings in a negative sign, and for consistency, one must require that both sides of the equation should vanish identically. Thus, the scalar field remains at $\phi(r)=\phi_{0}=0$, its asymptotic value and the resulting metric solution simply generates a Schwarzschild metric. This effectively rules out non-trivial scalar deformations of the Schwarzschild metric within a limit. The scalar no-hair theorem demonstrated is now a popular avenue of research [6, 7, 8, 9, 10]; however, there remains a curiosity if one can solve a system of static scalar field embedded within GR and weaken this constraint, even if slightly.
There are two further arguments which can provide strong impetus behind the curious case of a scalar field coupled to GR. The first argument is a motivation related to the study of cosmic expansion. In a spatially homogeneous cosmological universe, there is an omnipresent requirement of an entity beyond standard matter capable of accommodating the idea of an accelerated expansion [11, 12, 13, 14, 15]. A scalar field with a suitably chosen interaction profile has served as the most popular candidate in this pursuit [16, 17]. The present scope of simple scalar field cosmological models appears to be quite bleak in light of the steep fifth force constraints on the baryon-quintessence interaction profile [18]; nevertheless, scalar fields remain capable of generating solutions of considerable interest. As a specific example, one can cite the pseudo-Nambu-Goldstone-Boson (pNGB) model [19] as well as the models where a scalar field is capable of manipulating self-interaction via screening mechanisms [20, 21, 22].
@@
While one can consider $\lambda$ as a dimension-less parameter of the theory, $M(\phi)$ can lead to an evolving Higgs VEV $\nu$, calculated as
$\frac{\partial V}{\partial\phi}\bigg{\rvert}_{\nu}=0~{}~{},~{}~{}\nu=\sqrt{\frac{-2M(r)}{\lambda}}.$ (5)
It is to be noted that $\phi$ need not necessarily be the standard model Higgs. We can imagine the standard model Higgs $\varphi$ having two components, the Higgs particle $h$ alongwith a classical background field $\phi$, with a corresponding Lagrangian given by
@@
In this manuscript, we present a new exact spherically symmetric static solution of the Einstein field equation coupled with a self-interacting scalar field. The self-interaction of the scalar field is in the form written in Eq. 4. The static solution has no zero proper volume singularity at $r=0$. The curvature scalars are also non-singular for all coordinate ranges. We study the behavior of radial null geodesics for this solution and explain that depending on the regions of parameter space under consideration, the solution can describe a two-way or a one-way traversable Wormhole. There is no Schwarzschild limit to this solution. The scalar field behaves as a constant in the asymptotic limit and exhibits spontaneous symmetry breaking within the coordinate range associated with the formation of a wormhole throat. In other words, spontaneous symmetry breaking can be viewed as a threshold for the formation of a wormhole throat. Furthermore, we calculate the radius of the photon sphere, Lyapunov exponent, shadow radius, and innermost stable circular orbits for this spacetime and discuss their nature.
2 An Exact Solution and its Properties
@@
The scalar field $\phi$ is a function of radial coordinate $r$ alone. The self-interaction of the scalar field is taken in a form as in 4, i.e., $V(\phi)=V_{0}+\frac{1}{2}M(\phi)\phi^{2}+\frac{\lambda}{4}\phi^{4}$. We write a general static spherically symmetric metric tensor as
$ds^{2}=-G(r)dt^{2}+G(r)^{-1}dr^{2}+S(r)^{2}d\Omega^{2},$ (8)
where the metric coefficients are also functions of $r$ alone. The coordinates have natural domains, as in
@@
$G^{\prime\prime}-2G^{\prime}\sigma^{\prime}-V(\phi)=0,$ (11)
$G^{\prime\prime}-2G\left(2(\sigma^{\prime})^{2}-\sigma^{\prime\prime}\right)+2e^{2\sigma}=0.$ (12)
@@
It is straightforward to note that the scalars do not diverge for any value of $r$. In particular, there is no singularity at $r=0$ as long as $a$ is treated as a non-zero parameter and throughout this article we keep the parameter that way. The role of this parameter resembles the recently proposed solution of a regular black hole [39] and its time-evolving analogue [40]. We plot the scalars as a function of radial coordinate in 1 to prove our claim. For the purpose of this plot, we have chosen $C_{1}=1$ and $C_{2}=0.01$, while the parameter $a$ is being varied. Using the exact form of the metric coefficients, it is also straightforward to calculate the components of the energy-momentum tensor (using the mixed components of the field equations $G^{\mu}_{\nu}=8\pi G_{N}T^{\mu}_{\nu}$). The density is given by $\rho=-T^{t}_{t}$ and the radial pressure is given by $p_{r}=T^{r}_{r}$. Furthermore, we can check if there is any violation of Null Energy Condition (NEC) inside the spherical geometry. The necessary condition for an NEC to be valid is that the sum of density and radial pressure should be positive. Any violation of NEC is directly connected to the idea of a Null Convergence Condition (NCC), which is a characteristic of wormhole geometry, where radial null rays can defocus around the wormhole throat [41, 42, 43, 44, 45]. We plot the NEC profile in 2 as a function of $r$. For the graph on the top left of 2, we keep the two parameters $C_{1}$ and $C_{2}$ fixed at unity and vary the parameter $a$. For the graph on top right, we keep $a$ fixed at $0.5$ and $C_{2}$ fixed at $1$ while varying $C_{1}$ from $0.01$ to $1$. For the graph below, we keep $a$ fixed at $0.5$ and $C_{1}$ fixed at $1$ while varying $C_{2}$ from $0.1$ to $10$.
@@
We use 21 to check if there is any formation of the horizon for any value or range of $r$, for which we simply require that $\frac{dr}{dt}=0$ for any $r\in(-\infty,+\infty)$. Given the fact that 21 is not straightforward to solve, primarily due to the presence of a $\tan^{-1}$ term, we plot the RHS of the equation in 6. The graph on the top left is for a chosen set of $C_{1}=1$, $C_{2}=1$. We change the value of $a$ and find that as long as $a\lesssim 0.9$, the curve intersects zero, and we can, in principle, have a horizon at some coordinate location. However, this is definitely not a black hole since we do not have any curvature singularities. It can be rendered as a one-way wormhole with an extremal null throat. However, for $a\gtrsim 0.9$, we have $\frac{dr}{dt}\neq 0$ for any $r\in(-\infty,+\infty)$ and therefore the metric represents a two-way traversable wormhole. For this set, $a\sim 0.9$ defines a threshold of horizon formation, however, this is not an exhaustive analysis. For a different set $C_{1}=1$, $C_{2}=3$ (graph on top right), the threshold value of $a$ is found to be $\sim 1.25$. The bottom left graph is for a fixed set of $a=C_{2}=1$, while the parameter $C_{1}$ is varied, and for this set, we do not have any formation of horizon. The bottom right graph is for a fixed set of $a=C_{1}=1$, while the parameter $C_{2}$ is varied.
In order to have a mathematical argument behind this parameter-dependence of horizon formation we approximate 21 around a very small neighbourhood of $r=0$ (essentially to ignore any terms $\sim\textit{O}(r^{3})$) or higher. Under this approximation $\tan^{-1}(\frac{r}{a})\approx\frac{r}{a}$, and the radial null geodesic curve equation becomes a quadratic equation

### v3

# Traversable Wormholes with a Spontaneous Symmetry Breaking Scalar Field

@@ Abstract
We argue that a spherically symmetric traversable wormhole solution of the Einstein field equations can be supported by minimally coupled self-interacting scalar field which allows a spontaneous symmetry breaking of the field around the wormhole throat. We study two cases : (i) the phantom wormhole solution of Bronnikov and (ii) a generalized Kiselev wormhole. We study the property of radial null geodesics and show that the metric can describe either a two-way or a one-way traversable wormhole depending on certain parameter ranges. The scalar field exhibits spontaneous symmetry breaking within the coordinate range where a wormhole throat forms and helps one suggest that spontaneous symmetry breaking may act as a threshold for wormhole throat formation. We also compute the radius of the photon sphere, the Lyapunov exponent, the shadow radius, and the innermost stable circular orbits for the geometries.
@@ 1 Introduction
Although the General Theory of Relativity (GR) is widely regarded as the most accurate description of gravity to date, the number of exact solutions to its field equations remains relatively small. This limitation is primarily due to the challenges involved in solving the field equations: coupled partial differential equations that form the basis of this theory. Even under spherical symmetry the field equations of GR are inherently quite restrictive. A clear example of this rigidity is found in the fact that even for a vacuum (where $T_{\mu\nu}=0$), the Schwarzschild solution stands as the unique static and isotropic solution that satisfies the condition of asymptotic flatness [1, 2, 3]. This result is central to the formulation of Jebsen-Birkhoff’s theorem [4, 5]. The Schwarzschild solution represents a singularity with zero proper volume surrounded by a horizon. When matter is incorporated into the field equations, they become even more restrictive and less predictable. Some partial solutions have been obtained in specific cases, such as GR coupled to self-interacting scalar fields of dilaton or axion types or when coupled to Maxwell electrodynamics or a Yang-Mills field. However, static field configurations in GR are heavily constrained outside the event horizon, as demonstrated by the no-hair theorems.
A simple minimally coupled self-interacting scalar field $\phi$ contributes to the energy-momentum tensor through
@@
$V(\phi)$ is the self-interaction of the scalar field. The scalar field profile is imposed upon alongwith a requirement that in $r\rightarrow\infty$ limit the scalar field $\phi(r)$ should be a constant (one can choose to normalize this constant to zero). Moreover, a requirement of asymptotic flatness dictates that the self-interaction potential should have a local extremum at the origin of the scalar field. In fact, we can use these conditions to demonstrate the no-hair theorem starting from the scalar field evolution equation
@@
For a static case, only the $g^{rr}$ term contributes to the above equation. We multiply the equation with $\phi(r)$ and do integration by parts from the horizon to infinity, yielding
@@
Usually, one realizes a horizon $h$ at the largest zero of $g^{rr}$. Thus in principle it is possible to integrate the above equation for all $g^{rr}\geq 0$. If we assume regularity of $\phi$ and $\sqrt{-g}$ at the horizon, the boundary terms can be dropped (also keeping in mind an asymptotic fall-off of $\phi(r)$ as $r\rightarrow\infty$). In this context, the condition $\phi\frac{dV}{d\phi}\geq 0$ for all $r\geq h$ essentially implies that the self-interaction potential has a local minimum at $\phi=0$ and no local maximum in the $\phi$-range considered outside the horizon. Under this condition, the integrands on both sides are non-negative. However, the integral on LHS brings in a negative sign, and for consistency, both sides of the equation should vanish identically. Thus the scalar field remains at $\phi(r)=\phi_{0}=0$, its asymptotic value and the resulting metric solution simply generates a Schwarzschild metric. This effectively rules out non-trivial scalar deformations of the Schwarzschild metric within a limit. The scalar no-hair theorem demonstrated is now a popular avenue of research [6, 7, 8, 9, 10]; however, there remains a curiosity if one can solve a system of static scalar field embedded within GR and weaken this constraint, even if slightly.
There are two further arguments which provide strong impetus behind the curious case of a scalar field coupled to GR. The first argument is a motivation related to the study of cosmic expansion. In a spatially homogeneous cosmological universe, there is an omnipresent requirement of something beyond standard matter, capable of accommodating the idea of an accelerated expansion [11, 12, 13, 14, 15]. A scalar field with a suitably chosen interaction profile has served as the most popular candidate in this pursuit [16, 17]. The present scope of simple scalar field cosmological models appears to be bleak in light of the steep fifth force constraints on the baryon-quintessence interaction profile [18]; nevertheless, scalar fields remain capable of generating solutions of considerable interest. As a specific example, we refer to the pseudo-Nambu-Goldstone-Boson (pNGB) model [19] and models where a scalar field is capable of manipulating self-interaction via screening mechanisms [20, 21, 22].
@@
While one can consider $\lambda$ as a dimensionless parameter of the theory, the presence of $M(\phi)$ can lead to an evolving Higgs VEV $\nu$. It must be noted that $\phi$ need not be the standard model Higgs. We can imagine the standard model Higgs $\varphi$ having two components, the Higgs particle $h$ alongwith a classical background field $\phi$, with a corresponding Lagrangian given by
@@
In this article, we study the effect of introducing an $M(\phi)$ term in the scalar self-interaction $V(\phi)=V_{0}+M(\phi)\phi^{2}+\lambda\phi^{4}$ on two spherically symmetric static solutions of the Einstein field equations. The first metric tensor we discuss is a Phantom wormhole solution, initially derived by Bronnikov and Fabris for a minimally coupled scalar field with negative kinetic energy (sometimes favored cosmologically) [39]. We further prove that this wormhole metric can be rewritten in a generalized Kiselev form. Kiselev’s metric solution without any modification describes a black hole [40, 41]
$ds^{2}=-\left(1-{2m\over r}-{K\over r^{1+3w}}\right)dt^{2}+{dr^{2}\over 1-{2m\over r}-{K\over r^{1+3w}}}+r^{2}\,d\Omega_{2}^{2}.$ (6)
This model can describe three classes of solutions within one framework: Schwarzschild for $w=0$, Reissner-Nordstrom for $w=1/3$, and Schwarzschild-(anti)-de Sitter for $w=-1$. It has been proved that the Kiselev metric can not be supported by a perfect fluid or a simple quintessence. We regularize the metric tensor by replacing $r$ with $(r^{2}+a^{2})^{1/2}$ and make necessary adjustments in the other parameters to write it as
$g_{00}=1-\frac{2m}{\sqrt{a^{2}+r^{2}}}-\frac{p}{f(r)\left(a^{2}+r^{2}\right)^{\frac{1}{2}(3w+1)}}.$ (7)
We study the behavior of radial null geodesics for these two solutions and explain that depending on the regions of parameter space under consideration, the solutions behave as a traversable Wormhole. We show that it is possible to support these solutions using a scalar field which behaves as a constant in the asymptotic limit and exhibits spontaneous symmetry breaking (SSB) within the coordinate range associated with the formation of a wormhole throat. In other words, spontaneous symmetry breaking can be viewed as a threshold for the formation of a wormhole throat. Furthermore, we calculate the radius of the photon sphere, Lyapunov exponent, shadow radius, and innermost stable circular orbits for this spacetime and discuss their nature. The general requirement for non-trivial static solutions of the field equations of GR as well as of extensions of GR has always been steadfast, and there are quite a few solutions in literature (see for instance, [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]). Our construction can be applied to any of these solutions, and an appropriate association with SSB can be made.
2 The Phantom Wormhole: Solution and Properties
@@
The scalar field $\phi$ is a function of radial coordinate $r$ alone. The self-interaction of the scalar field is taken in a form as in 4, i.e., $V(\phi)=V_{0}+\frac{1}{2}M(\phi)\phi^{2}+\frac{\lambda}{4}\phi^{4}$. We write a general static spherically symmetric metric tensor
$ds^{2}=-G(r)dt^{2}+\frac{dr^{2}}{G(r)}+S(r)^{2}d\Omega^{2},$ (9)
Note, that $g_{11}=\frac{1}{g_{00}}=G(r)^{-1}$, therefore this metric is in a standard Schwarzschild form. The metric coefficients are functions of $r$ alone. The coordinates have natural domains, as in
@@
$G^{\prime\prime}-2G^{\prime}\sigma^{\prime}-V(\phi)=0,$
$G^{\prime\prime}-2G\left(2(\sigma^{\prime})^{2}-\sigma^{\prime\prime}\right)+2e^{2\sigma}=0.$
@@
It is straightforward to note that the scalars do not diverge for any value of $r$. In particular, there is no singularity at $r=0$ as long as $a$ is treated as a non-zero parameter and throughout this article we keep the parameter that way. The role of this parameter resembles the recently proposed solution of a regular black hole [57] and its time-evolving analogue [58]. We plot the scalars as a function of radial coordinate in 1 to prove our claim. For the purpose of this plot, we have chosen $C_{1}=1$ and $C_{2}=0.01$, while the parameter $a$ is being varied. We also note that these scalars do not go to zero at an asymptotic limit $r\rightarrow\infty$, as it is found for a standard Schwarzschild solution. For our metric tensor both of the scalars reach a non-zero constant value at this limit, which is anyway not unexpected given the functional form of $g_{00}=1+C_{1}+\frac{C_{2}}{2a^{3}}\left[ar+(a^{2}+r^{2})\tan^{-1}(r/a)\right]$. This is quite distinct in comparison to a Schwarzschild metric coefficient $\sim\left(1-2M/r\right)$, which can only go to $1$ in an asymptotic limit. For our solution, $g_{00}$ can reach a constant value (not equal to $1$) for all $a\neq 0$. In other words, the presence of the parameter $a$ encodes the departure from the standard Schwarzschild metric.
Using the exact form of the metric coefficients, it is also straightforward to calculate the components of the energy-momentum tensor (using the mixed components of the field equations $G^{\mu}_{\nu}=8\pi G_{N}T^{\mu}_{\nu}$). The density is given by $\rho=-T^{t}_{t}$ and the radial pressure is given by $p_{r}=T^{r}_{r}$. Furthermore, we can check if there is any violation of Null Energy Condition (NEC) inside the spherical geometry. The necessary condition for an NEC to be valid is that the sum of density and radial pressure should be positive. Any violation of NEC is directly connected to the idea of a Null Convergence Condition (NCC), which is a characteristic of wormhole geometry, where radial null rays can defocus around the wormhole throat [59, 60, 61, 62, 63]. We plot the NEC profile in 2 as a function of $r$. For the graph on the top left of 2, we keep the two parameters $C_{1}$ and $C_{2}$ fixed at unity and vary the parameter $a$. For the graph on top right, we keep $a$ fixed at $0.5$ and $C_{2}$ fixed at $1$ while varying $C_{1}$ from $0.01$ to $1$. For the graph below, we keep $a$ fixed at $0.5$ and $C_{1}$ fixed at $1$ while varying $C_{2}$ from $0.1$ to $10$.
@@
We use 22 to check if there is any formation of the horizon for any value or range of $r$, for which we simply require that $\frac{dr}{dt}=0$ for any $r\in(-\infty,+\infty)$. Given the fact that 22 is not straightforward to solve, primarily due to the presence of a $\tan^{-1}$ term, we plot the RHS of the equation in 6. The graph on the top left is for a chosen set of $C_{1}=1$, $C_{2}=1$. We change the value of $a$ and find that as long as $a\lesssim 0.9$, the curve intersects zero, and we can, in principle, have a horizon at some coordinate location. However, this is definitely not a black hole since we do not have any curvature singularities. It can be rendered as a one-way wormhole with an extremal null throat. However, for $a\gtrsim 0.9$, we have $\frac{dr}{dt}\neq 0$ for any $r\in(-\infty,+\infty)$ and therefore the metric represents a two-way traversable wormhole. For this set, $a\sim 0.9$ defines a threshold of horizon formation, however, this is not an exhaustive analysis. For a different set $C_{1}=1$, $C_{2}=3$ (graph on top right), the threshold value of $a$ is found to be $\sim 1.25$. The bottom left graph is for a fixed set of $a=C_{2}=1$, while the parameter $C_{1}$ is varied, and for this set, we do not have any formation of horizon. The bottom right graph is for a fixed set of $a=C_{1}=1$, while the parameter $C_{2}$ is varied.
In order to have a mathematical argument behind this parameter-dependence of horizon formation, we approximate 22 around a very small neighborhood of $r=0$ (essentially to ignore any terms $\sim\textit{O}(r^{3})$) or higher. Under this approximation $\tan^{-1}(\frac{r}{a})\approx\frac{r}{a}$, and the radial null geodesic curve equation becomes a quadratic equation
@@
It can be proved that the parameter $a$ is nothing but the coordinate value at which a wormhole throat forms. This is evident through a straightforward coordinate transformation, $r^{2}+a^{2}=l^{2}$, applied to the metric in 9. The transformation leads to a domain for the radial coordinate $l\in(a,+\infty)$, indicating that now $a$ can be treated as the minimum value of the radial coordinate. As a consequence, the metric tensor in terms of $l$ takes on a new form
$ds^{2}=-G_{1}(l)dt^{2}+\frac{dr^{2}}{\left(1-\frac{a^{2}}{l^{2}}\right)G_{1}(l)}+l^{2}d\Omega^{2},$ (25)
In this framework, the modified $g_{11}$ term resembles a wormhole throat function [58]. A wormhole throat is defined as a closed, space-like two-dimensional surface with minimal area. To analyze this, one has to examine an embedding geometry of the metric on a constant time slice, where $\theta=\pi/2$. This leads to
$d\sigma^{2}=B^{2}dl^{2}+l^{2}d\phi^{2}~{}~{};~{}~{}B^{2}=\frac{1}{\left(1-\frac{a^{2}}{l^{2}}\right)G_{1}(l)}.$ (26)
$d\sigma^{2}$ can be mapped to a generic cylindrically symmetric metric in $3$-dimension
$d\sigma^{2}=dz^{2}+d\rho^{2}+\rho^{2}d\phi^{2},$ (27)
defined on the surface of revolution $\rho(z)$. Comparing these two equations we establish the correlation
$\rho^{2}=l^{2}~{};~{}dz^{2}+d\rho^{2}=B^{2}dl^{2},$ (28)
leading further to the conditions
$\frac{d\rho}{dz}=\frac{1}{(B^{2}-1)^{\frac{1}{2}}}~{}~{};~{}~{}\frac{d^{2}\rho}{dz^{2}}=-\frac{\frac{dB}{dl}B}{(B^{2}-1)^{2}}.$ (29)
In a cylindrically symmetric coordinate system, the throat of a wormhole can be visualized in an embedding diagram as a circle of radius $\rho$ on a surface of revolution. At the throat, the radius $\rho(z)$ should reach its minimum. A straightforward calculation can be done to show that in the limit $l\rightarrow a$, $\rho(z)$ indeed attains its minimum, satisfying the following two conditions
$\frac{d\rho}{dz}\Big{|}_{l\rightarrow a}=0~{}~{};~{}~{}\frac{d^{2}\rho}{dz^{2}}\Big{|}_{l\rightarrow a}>0.$ (30)
3 The Generalized Kiselev Wormhole: Solution and Properties
Using a simple algebraic manipulation one can rewrite the Phantom wormhole metric in 15 as
$G(r)=1-\frac{2m}{\sqrt{a^{2}+r^{2}}}-\frac{p(r)}{\left(a^{2}+r^{2}\right)^{\frac{1}{2}(3w+1)}},$ (31)
where
$p(r)=C_{1}\left(a^{2}+r^{2}\right)^{\frac{3(w+1)}{2}}-2m\left(a^{2}+r^{2}\right)^{\frac{3w}{2}}+\frac{C_{2}\left(a^{2}+r^{2}\right)^{\frac{3w}{2}+\frac{1}{2}}\left\{a^{2}\tan^{-1}\left(\frac{r}{a}\right)+r^{2}\tan^{-1}\left(\frac{r}{a}\right)+ar\right\}}{2a^{3}}.$ (32)
This is nothing but a generalization of the Kiselev metric [40] for which the function $p(r)$ is a number. Note that the parameter $w$ remains unperturbed, which, in the standard Kiselev metric, allows one to toggle between three classes of solutions, namely, Schwarzschild for $w=0$, Reissner-Nordstrom for $w=1/3$ and Schwarzschild-(anti)-de Sitter for $w=-1$. Since there seems to be a way to associate an SSB-breaking scalar field with the Phantom wormhole, it is a natural curiosity to ask if the same can also be done for a Kiselev metric or its generalization. We propose the following generalization
$G(r)=1-\frac{2m}{\sqrt{a^{2}+r^{2}}}-\frac{p}{f(r)\left(a^{2}+r^{2}\right)^{\frac{n(r)}{2}}},$ (33)
where
$\sigma(r)=-\frac{1}{2}\ln(r^{2}+a^{2})~{}~{},~{}~{}\phi(r)=C_{3}\pm 2\tan^{-1}(r/a)~{}~{},~{}~{}n(r)=(3w+1),$
$f(r)=a^{-2}p^{-1}-a^{2}C_{1}pr^{2}\left(a^{2}+r^{2}\right)^{\frac{3w}{2}+\frac{1}{2}}+a^{2}C_{2}pr^{2}\left(a^{2}+r^{2}\right)^{\frac{3w}{2}+\frac{1}{2}}\tan^{-1}\left(\frac{r}{a}\right)-2a^{2}m\left(a^{2}+r^{2}\right)^{\frac{3w}{2}}-$
$mr^{2}\sqrt{\frac{a^{2}+r^{2}}{a^{2}}}\left(a^{2}+r^{2}\right)^{\frac{3w}{2}}-a^{2}m\sqrt{\frac{a^{2}+r^{2}}{a^{2}}}\left(a^{2}+r^{2}\right)^{\frac{3w}{2}}-a^{4}C_{1}p\left(a^{2}+r^{2}\right)^{\frac{3w}{2}+\frac{1}{2}}$
$+a^{4}C_{2}p\left(a^{2}+r^{2}\right)^{\frac{3w}{2}+\frac{1}{2}}\tan^{-1}\left(\frac{r}{a}\right)+a^{3}C_{2}pr\left(a^{2}+r^{2}\right)^{\frac{3w}{2}+\frac{1}{2}}.$
Figure 7: Ricci scalar $R(r)$ and Kretschmann scalar $K(r)$ as a function of $r$ for different values of $a$ for a generalized Kiselev metric
The above set of expressions can be found by solving the field equations defined in 13 alongwith the metric ansatz 33. The first thing we check and confirm is that the curvature scalars, namely the Ricci and Kretschmann scalar do not diverge for any value of $r$. In particular, there is no singularity at $r=0$ as long as $a$ is treated as a non-zero parameter. Once again this parameter can be treated as a wormhole throat radius. We plot the scalars as a function of radial coordinate in 7, choosing particular values of $C_{1}$, $C_{2}$, $m$ and $p$, varying the parameter $a$ alone. The qualitative behavior of the scalars remains similar even for different values of the parameters. We note that these scalars become constant at an asymptotic limit $r\rightarrow\infty$, however, this constant can be scaled to zero for an appropriately chosen set of parameter values.
$R\Rightarrow\Big{(}-\frac{2m}{a^{3}}-2C_{1}p+\pi C_{2}p\Big{)},$
$K\Rightarrow\frac{60p(2C_{1}-\pi C_{2})}{a^{3}}+\frac{4}{a^{6}}\Bigg{\{}\frac{98}{\Big{(}a^{3}p(2C_{1}-\pi C_{2})+2\Big{)}^{2}}-\frac{2}{a^{3}p(2C_{1}-\pi C_{2})+4}+\frac{1}{\Big{(}a^{3}p(2C_{1}-\pi C_{2})+4\Big{)}^{2}}$
$-\frac{16}{a^{3}p(2C_{1}-\pi C_{2})+6}+\frac{16}{\Big{(}a^{3}p(2C_{1}-\pi C_{2})+6\Big{)}^{2}}-\frac{54}{a^{3}p(2C_{1}-\pi C2)+8}+\frac{81}{\Big{(}a^{3}p(2C1-\pi C2)+8\Big{)}^{2}}$
$+\frac{72}{a^{3}p(2C_{1}-\pi C_{2})+2}+73\Big{\}}+6p^{2}(\pi C_{2}-2C_{1})^{2}.$
One can calculate the components of the energy-momentum tensor for this metric using the field equations $G^{\mu}_{\nu}=8\pi G_{N}T^{\mu}_{\nu}$. We study, in particular, the NEC defined as $-T^{t}_{t}+T^{r}_{r}$ and plot the profile in 8 as a function of $r$. The graph on the left is for positive values of $C_{2}$, and the one on the right is for negative $C_{2}$. It is clearly seen that the NEC is violated at least in the present universe, i.e., in the coordinate range $0<r<\infty$.
Figure 8: Null Energy Condition as a function of $r$ for different values of $C_{2}$. $C_{1}=C_{2}=1$, $a=0.5$. Evolution for positive and negative values of $C_{2}$ are shown on left and right.
Figure 9: Plot of $M(\phi)$ as a function of $r$ for $\phi(r)=C_{3}+2\tan^{-1}(r/a)$. Top left: different values of $m$ are considered while the other parameters are fixed. Top right: different values of $p$ are considered while the other parameters are fixed. Bottom left: different values of $C_{3}$ are considered while the other parameters are fixed. Bottom right : evolution for $C_{3}>>0$.
We note once again that the solutions discussed in the article are consistent if and only if the term $M(\phi)$ in the self-interaction potential is a function. Using 13 and 33, we solve for $M(\phi)$, and in 9, we plot it as a function of $r$. The graph on the top left is for different values of $a$ while the three other parameters $C_{1}$, $C_{2}$ and $C_{3}$ are fixed. The graph on the top left shows $M(r)$ for different values of $m$ while the other parameters are fixed. The graph on top right shows $M(r)$ for different values of $p$ while the other parameters are fixed. The bottom left graph shows $M(r)$ profile for different values of $C_{3}$. There is always a switch from negative into positive values of $M(\phi)$ within a small neighborhood of $r=0$, but never exactly at $r=0$, signifying the spontaneous breakdown of $Z_{2}$ symmetry. However, unlike a Phantom wormhole metric, it is possible in this case to have a solution with no SSB. This can be seen by taking values of $C_{3}>>C2$. The $M(r)$ profile corresponding to this case is shown in the bottom right graph of 9.
Figure 10: $\frac{dr}{dt}$ as a function of $r$ for different set of initial conditions. Top left: The value of $a$ is varied while the other parameters are fixed. Top right: The value of $C_{2}$ is varied while the other parameters are fixed. Bottom left: The parameter $m$ is varied while the other parameters are fixed. Bottom right: The parameter $p$ is varied while the other parameters are fixed.
We study the evolution of $g_{00}$ component of the generalized Kiselev metric to see if $\frac{dr}{dt}=0$ for any value of $r\in(-\infty,+\infty)$, which marks the formation of a horizon. We plot $\frac{dr}{dt}$ as a function of $r$ in 10. It is straightforward to deduce that there are no zeros of $\ \frac{dr}{dt}$ for most of the parameter space. The metric tensor describes a two-way traversable wormhole apart from a few exceptions where one might have a formation of the horizon without any central singularity at $r=0$ or a regular black hole.